Students captured 10 temperature readings in Celsius on a particular day. Write an Algorithm to input ten values and get the average temperature for that day. if the average temperature value is in between 970 Fahrenheit and 990 Fahrenheit then display the message “Your body temperature is normal…”. If it is more than 100.40Fahrenheit then display the message “You have a fever caused by an infection or illness…


1
Expert's answer
2021-12-02T07:31:39-0500

input:

array tc(10) \\ 10 temperature readings in Celsius


sum=0

for i from 1 to 10

sum=sum+tc[i]


sc=sum/10 \\ average temperature in Celsius

sf=sc*1.8+32 \\ average temperature in Fahrenheit

if 97<=sf<=99 then

display: “Your body temperature is normal…”

else if sf>100.4 then

display: “You have a fever caused by an infection or illness"